OpenStudy (anonymous):

Devon and his friends bought strawberry wafers for $3 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $30 to buy a total of 22 packets of wafers of the two varieties. Part A: Write a system of equations that can be solved to find the number of packets of strawberry wafers and the number of packets of chocolate wafers that Devon and his friends bought at the carnival. Define the variables used in the equations. (5 points)

OpenStudy (anonymous):

Part B: How many packets of chocolate wafers and strawberry wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer. (5 points)

OpenStudy (ddcamp):

Write one equation that finds the total number of wafers (if they bought s strawberry wafers and c chocolate wafers, how many wafers did they buy total?) The other equation should be the amount they spent (i.e. If they bought s strawberry wafers and c chocolate wafers, how much did they spend?)

OpenStudy (anonymous):

Ok I guess.

OpenStudy (ddcamp):

I actually have to go soon, so here's sort of what the equations look like, hope they help. \[s + c = total\ wafers \\ 3 \times s + 1 \times c = total\ cost\]

OpenStudy (anonymous):

Ok, thanks.